# MFAS: Emotion Recognition through Multiple Perspectives Fusion Architecture Search Emulating Human Cognition

### v2

@@ Abstract
Speech Emotion Recognition (SER) is an important research topic in human-computer interaction. Many recent works directly extract emotional cues using pre-trained knowledge, often disregarding the appropriateness of it. Taking inspiration from human cognition, we partition speech information into Textual-related Emotional Content (TEC), approximating text, and Speech-related Emotional Content (SEC), encapsulating acoustic information. Given that humans perceive TEC and SEC as discrete and dimensional respectively, we capture them using quantization-based and continuous-based knowledge respectively. Moreover, considering the diverse contextual demands for adjusting TEC using different aspects of SEC as humans do, we propose a new search space for SER and employ neural architecture search for automated design of optimal adjustment strategy. Building on these results, we propose a novel SER framework named Multiple perspective Fusion Architecture Search (MFAS). Experimental results show that MFAS outperforms existing methods.
*Equal contribution
Index Terms—  speech emotion recognition, multiple perspectives, neural architecture search, different modeling methods.
@@ 1 Introduction
Due to its significant contribution to human-computer interaction, Speech Emotion Recognition (SER) has received increasing attention. Researchers have been working towards endowing models with the ability to perceive and recognize emotions akin to humans. While some studies have incorporated various prior knowledge to guide the modeling process, the emergence of highly performant pre-trained models has led many recent works to directly leverage pre-trained knowledge for favorable outcomes. Nevertheless, the absence of prior knowledge in modeling could potentially restrict the model’s ability to extract emotional cues from data. In this paper, we employ pre-trained knowledge and simulate human cognitive processes to model emotional cues in speech, and we validate the significance of such biased induction in SER.
Fig. 1: The overall framework of MFAS.
While both text and speech are vital for emotion recognition [1, 2, 3], obtaining accurate transcribed text in real-world scenarios can be challenging. Thus, we broadly partition emotional cues in speech into two types: Textual-related Emotional Content (TEC), which can be extracted using Automated Speech Recognition (ASR) model to approximate text, and Speech-related Emotional Content (SEC), which can be extracted using pre-trained models to emphasize acoustic information. Many existing pre-trained ASR models, grounded in inductive bias towards pronunciation units, introduce quantization operations to enhance the model’s perception of text [4, 5, 6]. After undergoing two stages of pre-training and fine-tuning, they can efficiently construct a constrained set of discrete units. This discrete knowledge aligns well with the cognitive process of human understanding of TEC in speech.
Unlike TEC, during SEC extraction, existing methods lack explicit biased induction guidance and often rely on leveraging knowledge from pre-trained models [7, 8]. However, emotional information perceived by humans is dimensional, continuous [9, 10]. We follow this cognitive principle and compare the performance of quantitative modeling pre-trained knowledge [5] with continuous modeling pre-trained knowledge [11] in the SER fairly. Experimental results show that pre-trained knowledge based on continuous modeling is better suited for capturing SEC.
The proper utilization of multi-perspective information, such as TEC and SEC, is also a matter of concern. People typically perceive emotions based on explicit emotional words in text. However, in certain contexts, this information can be misleading, such as someone might say ”I’m really happy right now” with emphasis when they are angry. In such cases, individuals adjust their perception of emotional words based on acoustic information like intonation or tone, or both. We follow this principle and utilize SEC to adjust the model’s perception of TEC. However, due to the varying significance of multi-level speech information [12], selecting the optimal level of information is imperative. While exhaustive selection can yield answers, it’s a time-consuming and resource-intensive process. Therefore, we propose a novel search space for SER and utilize Neural Architecture Search (NAS) [13] to automatically design adjustment strategy. Experimental results analyze the necessity of this process.
Building on the above findings, we propose a novel SER framework, called Multiple perspective Fusion Architecture Search (MFAS). As shown in Figure 1, MFAS emulates human cognition by employing diverse knowledge perspectives to extract TEC and SEC from speech, and also offers the capability to automatically utilize SEC to adjust biases in TEC. Finally, both features are employed for emotion analysis. The contributions of this paper can be summarized as follows:
  • We validate that knowledge learned by the pre-trained method based on continuous modeling is better suited for capturing SEC than quantization modeling.
  • We propose a novel search space for SER, and utilize NAS to design adjustment strategy for TEC using SEC.
  • We propose a novel SER framework, called MFAS, to simulate the human cognition. Experimental results show that MFAS outperforms existing methods.
Fig. 2: Unified pre-training framework for different modeling methods. Simultaneously, testing the ability of pre-trained knowledge to capture emotional cues.
@@ 2 Datasets
Interactive Emotional Dyadic Motion Capture (IEMOCAP) [14] dataset is a widely adopted benchmark for evaluating emotion recognition models. Following prior research, we merge the ”happy” and ”excited” categories into a single ”happy” and focus on four emotions: angry, sad, happy, and neutral. To ensure a robust evaluation of our method, we adopted the 5-fold leave-one-session-out and 10-fold leave-one-speaker-out validation strategies to generate results.
@@
3 Experimental Setup
@@
Additionally, we employed an SGD optimizer with a learning rate of 5 to optimize the operation weight $\alpha$, as described in Section 5.1. When dealing with MFAS dual-stream inputs, we double the dimension of the encoding hidden features for both spectrograms and MFCCs. These features are evenly distributed to each stream for co-attention calculations.
@@ V A D UA(%) WA(%) V A D UA(%) WA(%) UA(%) WA(%)
4.1 Description of different modeling methods
Currently, speech modeling methods can be broadly categorized into two approaches: quantization-based reconstruction [4, 5, 6] and continuous frame-based reconstruction [16, 11]. Both approaches aim to understand the content of speech by reconstructing the masked portions. The key difference between these approaches lies in their representation space of speech frames. Formally, we represent the speech as $X=[x_{1},...,x_{T}]$, where $T$ is the number of frames. Given a set of indices $Z=[z_{1},...,z_{T^{\prime}}]$ corresponding to the frames that need to be masked, where $0\leq T^{\prime}\leq T$, the frames at the indices in $Z$ are replaced with a trainable code word $c$, resulting in $\tilde{X}=[...,x_{t},...,c,...]$, where $0\leq t<T$. The model then encodes $\tilde{X}$ to obtain its output $\hat{X}=[\hat{x}_{1},...,\hat{x}_{T}]$.
4.2 Hyperparameter Configuration
As shown in Figure 2, we use the same model architecture for SER under different modeling constraints to fairly compare them. The specific parameters of the model structure are shown in Table 1. Following previous works [5, 11], we first utilize a 1D-CNN network to encode the speech and obtain $X$. In the quantization modeling approach, we imitate Wav2vec2 [5] and assign $B$ codebooks to all speech frames, each containing $W$ codewords, resulting in a total of $N=W^{B}$ clustering centers. We uniformly set the mask rate to 65%, and the label vector $Y$ is composed of the clustering centers corresponding to $X[Z]$ and the sampled negative examples. The model is trained with BCEloss for reconstruction. On the other hand, in the continuous modeling approach, we imitate Data2vec [11] and use the mean value of the outputs of the last $L$ encoder layers as the $Y$. The model is trained with MSEloss for reconstruction.
4.3 Experimental results
To comprehensively compare the capabilities of various modeling approaches in capturing SEC, we not only test the models’ capability in discrete emotion recognition but also in dimensional emotion analysis. To highlight the differences between regression results, we utilize mean squared error (MSE) as the evaluation metric for dimensional emotion analysis. In Table 2, we report the MSE metrics of valence (V), activation (A), dominance (D), and report unweighted accuracy (UA), weighted accuracy (WA).
As demonstrated by the performance in V, A, and D, in dimensional emotion analysis, knowledge acquired through quantization modeling exhibits notable limitations. The performance significantly improves as the number of quantization units, $N$, increases. However, it is only when the number of $N$ is sufficiently large that this method approaches the performance of the continuous modeling. In the continuous modeling, the variation of the encoding layers $L$ as reconstructed labels does not noticeably affect performance.
@@ 4.3 Experimental results
Based on the above experiments, we can conclude that leveraging knowledge acquired through pre-training methods based on continuous modeling enables the capture of more comprehensive emotional information from speech. Whether in discrete emotion recognition or dimensional emotion analysis, the pre-trained knowledge derived from continuous modeling proves to be a superior choice for extracting SEC.
5 Multiple Perspective Fusion Architecture Search
Table 3: Performance comparison with existing works. Evaluation measures are UA / WA for IEMOCAP and UAR / WAR for SAVEE
Model Year
IEMOCAP
Leave-one-session
Model Year
IEMOCAP
Leave-one-speaker
Model Year SAVEE
IS09-classification [17] 2019 63.8 / 68.1 CNN TF Att.pooling [18] 2018 68.1 / 71.8 DCNN [19] 2020    -    / 82.1
CME [20] 2021 73.5 / 72.7 HNSD [21] 2021 72.5 / 70.5 TSP+INCA [22] 2021 83.4 / 84.8
Co-attention [7] 2022 71.1 / 69.8 Co-attention [7] 2022 72.7 / 71.6 CPAC [23] 2022 83.7 / 85.6
Prosody2Vec [24] 2022 73.3 / 72.4 Prosody2Vec [24] 2022 73.9 / 72.7 TIM-Net [25] 2023 77.3 / 79.4
MFAS 2023 74.0 / 71.9 MFAS 2023 74.6 / 73.2 MFAS 2023 86.0 / 86.3
5.1 Description of search space
As illustrated in Figure 1, utilizing an ASR pre-trained model based on quantization modeling allows us to obtain Textual-related Emotional Content (TEC), referred to as $X^{t}$. Similarly, employing a pre-trained model based on continuous modeling with $k$ encoding layers enables us to acquire different levels of Speech-related Emotional Content (SEC), denoted as $[X_{1}^{s},...,X_{m}^{s},...,X_{k}^{s}]$, where $m=\frac{k}{2}$. As the intermediate information generated during speech understanding carries different information [26], we regard $X_{1}^{s}$, $X_{m}^{s}$, and $X_{k}^{s}$ as three levels of information.
  • To extract $X_{k}^{s}$ from speech, the intermediate layer need to capture deeper semantic information to fully understand content, denoted as $X_{m}^{s}$.
To prevent interference among different levels of SEC and effectively train the model to adjust TEC using it, during each mini-batch training, we randomly sample one level of SEC, $X^{s}_{i}$, where $i\in[1,m,k]$, for calculation. To ensure that $X^{s}_{1}$ and $X^{s}_{m}$ do not contain target-related information, we detach $X^{s}_{i}$ before using it. We offer a total of eight adjustment operations in the operation pool $\mathbb{O}$, and each operation takes two tensor inputs and yields a tensor output:
$\small\mathrm{Zero}(X^{s}_{i},X^{t})=0$ (1)
$\small\mathrm{Sum}(X^{s}_{i},X^{t})=X^{s}_{i}+X^{t}$ (2)
$\mathrm{ConcatFC}(X^{s}_{i},$ $X^{t})=$ (4)
$\mathrm{ReLU}($ $\mathrm{Linear}(\mathrm{Concat}(X^{s}_{i},X^{t})))$
$\mathrm{where}\ H=\mathrm{Linear}(X^{t})$
Attention, ConcatFC, and ISM operations have corresponding reverse versions. For example, Attention_r$(X^{s}_{i},X^{t})$ = Attention$(X^{t},X^{s}_{i})$.
Following that, within the Fusion Cell, we employ a differentiable search algorithm [13] to design the adjustment operations. The computational process can be described as:
where $\alpha\in\mathbb{R}^{|\mathbb{O}|}$ is a trainable weight that measures the importance of different operations, and $|\mathbb{O}|$ denotes the number of operations in $\mathbb{O}$. Both $X^{s}_{k}$ and $X^{t}_{f}$ are subsequently used for emotional analysis as SEC and TEC, respectively.
5.2 Experimental results
5.2.1 Comparison with existing works
Following the conclusions in Section 4, in the experiments of this chapter, we employ the pre-trained knowledge based on continuous modeling from Data2vec to extract SEC from speech. For TEC, we retain the prior knowledge of pronunciation units and capture it using fine-tuned Wav2vec2. Following prior works, we conducted performance tests of MFAS on both the IEMOCAP and SAVEE datasets. As depicted in Table 3, in our experiments on the IEMOCAP dataset, MFAS attains the highest UA score in the Leave-one-session cross-validation strategy, demonstrating competitive WA scores even in the presence of class imbalance. Furthermore, within the Leave-one-speaker cross-validation strategy, MFAS consistently achieves the top UA and WA scores. When applied to the broader range of categories in the SAVEE dataset, MFAS exhibits a significant advantage, surpassing the sota in the 10-fold cross-validation.
@@ 5.2.1 Comparison with existing works
5.2.2 Comparison with different configurations
We analyze the importance of each module in MFAS by evaluating various model configurations using the Leave-one-session strategy on IEMOCAP. As shown in Table 4, consistent with the findings in Section 4, the pre-trained knowledge based on continuous modeling in Data2vec better captures emotional cues in speech. However, simply incorporating TEC information without adjustment introduces bias, resulting in no significant performance improvement and even a slight decrease in the WA metric. Through the adjustment strategy search by MFAS, the model leverages SEC to mitigate bias in TEC, leading to a further improvement in model performance.
5.2.3 Visualization of Search Results
As shown in Figure 3, in various data partition scenarios, the model automatically selects the optimal level of SEC to adjust TEC. To further demonstrate the performance of MFAS on imbalanced datasets, we present the performance confusion matrix of MFAS compared to the Co-attention method using only Wav2vec2 as the TEC Encoder, as shown in Figure 4.
Fig. 3: The visualization of adjustment strategy search results for the Leave-one-session strategy. Here, $1$, $m$, and $k$ correspond to three levels of information. The red color indicates the selected paths.

### v3

# MFSN: Multi-perspective Fusion Search Network For Pre-training Knowledge in Speech Emotion Recognition

@@ Abstract
Speech Emotion Recognition (SER) is an important research topic in human-computer interaction. Many recent works focus on directly extracting emotional cues through pre-trained knowledge, frequently overlooking considerations of appropriateness and comprehensiveness. Therefore, we propose a novel framework for pre-training knowledge in SER, called Multi-perspective Fusion Search Network (MFSN). Considering comprehensiveness, we partition speech knowledge into Textual-related Emotional Content (TEC) and Speech-related Emotional Content (SEC), capturing cues from both semantic and acoustic perspectives, and we design a new architecture search space to fully leverage them. Considering appropriateness, we verify the efficacy of different modeling approaches in capturing SEC and fills the gap in current research. Experimental results on multiple datasets demonstrate the superiority of MFSN.
keywords:
speech emotion recognition, multiple perspectives, neural architecture search
*Equal contribution; **Corresponding authors
@@ 1 Introduction
Due to its significant contribution to human-computer interaction, Speech Emotion Recognition (SER) has received increasing attention. Researchers have been working towards endowing models with the ability to perceive and recognize emotions akin to humans. While some studies have incorporated various prior knowledge to guide the modeling process, the emergence of highly performant pre-trained models has led many recent works to directly leverage pre-trained knowledge for favorable outcomes. However, due to differences in training objectives, pre-training modeling methods and knowledge that excel in other downstream tasks may not perform well in SER. In this paper, we undertake a consideration and analysis from both the perspectives of appropriateness and comprehensiveness, proposing a pre-training knowledge utilization framework tailored for SER.
From the perspective of comprehensiveness, both text and speech play pivotal roles in emotion recognition [1, 2, 3]. However, obtaining accurate transcriptions in real-world scenarios poses challenges, and directly utilizing semantic information in speech does not enable its comprehensive exploration. Consequently, we categorize emotional cues in speech into two types: Textual-related Emotional Content (TEC), which can be extracted using Automated Speech Recognition (ASR) model to approximate text, and Speech-related Emotional Content (SEC), which can be extracted using pre-trained model to represent acoustic feature.
Figure 1: The overall framework of MFSN.
Directly leveraging both forms of information proves impractical, as transcriptions often contains emotionally charged words that may deviate from the true emotional context, potentially causing misinterpretations. For instance, someone might emphatically say, ”I’m really happy right now” when they are actually angry. In such instances, individuals recalibrate their perception of emotional words based on acoustic information such as intonation or tone, or others. We follow this principle and leverage speech feature to refine the model’s understanding of TEC. In light of the varying significance of multi-level speech information [4], the thoughtful selection of the optimal level, along with the fusion operation with TEC, becomes crucial. While exhaustive selection may yield insights, it is a time-consuming and resource-intensive process. Hence, we propose the use of Neural Architecture Search (NAS) for the efficient accomplishment of the aforementioned tasks. Specifically, we design a noval search space and employ a differentiable search algorithm [5] for automated strategy design.
From the perspective of appropriateness, many contemporary approaches incorporate prior knowledge to tailor modeling methods for the target task. For instance, many existing pre-trained ASR models, grounded in inductive bias towards pronunciation units, introduce quantization operations to enhance the model’s perception of text [6, 7, 8]. Following two stages of pre-training and fine-tuning, these models can construct a constrained set of discrete units. However, there is currently no consensus on the modeling methods for SEC, often relying on pre-training speech models without due consideration [9, 10]. Given that human emotions are often dimensional and continuous [11, 12], we follow this principle and systematically compare the performance of quantitative modeling pre-trained knowledge [7] with continuous modeling pre-trained knowledge [13] in SER. This not only offers insights into the modeling methods for SEC but also fills current research gaps.
In conclusion, considering both comprehensiveness and appropriateness, we propose a novel framework for pre-training knowledge in SER, called Multi-perspective Fusion Search Network (MFSN). Illustrated in Figure 1, MFSN employs continuous-based knowledge to capture SEC and utilizes quantization-based knowledge to capture TEC. Furthermore, it possesses the capability to automatically leverage the optimal level speech feature to refine the understanding of TEC. Finally, both features are employed for emotion analysis. The contributions of this paper can be summarized as follows:
  • We propose a novel framework for using pre-trained knowledge in SER, named MFSN. MFSN extract both TEC and SEC from appropriate perspectives. It automatically designs adjustment strategies for TEC within a newly designed search space, comprehensively exploring emotional cues.
  • For the first time from a modeling perspective, we verify that pre-training methods based on continuous modeling are more suitable for capturing SEC than quantization.
  • Experimental results on multiple SER benchmark datasets demonstrate that the performance of MFSN achieves state-of-the-art levels.
Table 1: Performance comparison with existing works. Evaluation measures are UA / WA for IEMOCAP and UAR / WAR for SAVEE.
Model
IEMOCAP
Leave-one-session
Model
IEMOCAP
Leave-one-speaker
Model SAVEE
IS09-classification[14] 63.8 / 68.1 CNN TF Att.pooling[15] 68.1 / 71.8 DCNN[16]    -    / 82.1
CME[17] 73.5 / 72.7 HNSD[18] 72.5 / 70.5 TSP+INCA[19] 83.4 / 84.8
Co-attention[9] 71.1 / 69.8 Co-attention[9] 72.7 / 71.6 CPAC[20] 83.7 / 85.6
Prosody2Vec[21] 73.3 / 72.4 Prosody2Vec[21] 73.9 / 72.7 TIM-Net[22] 77.3 / 79.4
MFSN 74.0 / 71.9 MFSN 74.6 / 73.2 MFSN 86.0 / 86.3
2 Multiple-perspective Fusion Search Network
2.1 TEC & SEC
In order to comprehensively extract emotional cues from speech, we partition the emotion-related content in speech into two types: Textual-related Emotional Content (TEC) and Speech-related Emotional Content (SEC). As illustrated in Figure 1, utilizing an ASR pre-trained model based on quantization modeling allows us to obtain TEC, referred to as $X^{t}$. Similarly, employing a pre-trained model based on continuous modeling with $k$ encoding layers enables us to acquire different levels of speech features, denoted as $[X_{1}^{s},...,X_{m}^{s},...,X_{k}^{s}]$, where $m=\frac{k}{2}$, $X_{k}^{s}$ is SEC. As the intermediate information generated during speech understanding carries different information [23], we regard $X_{1}^{s}$, $X_{m}^{s}$, and $X_{k}^{s}$ as three levels of information.
  • To extract $X_{k}^{s}$ from speech, the intermediate layer need to capture deeper information to fully understand content, denoted as $X_{m}^{s}$.
2.2 Search Space
Figure 2: The search space we designed is divided into two parts: Choice Cell and Fusion Cell.
As illustrated in Figure 2, to efficiently accomplish the selection of multi-level features and fusion operations, we design a novel search space. Firstly, to endow the model with the capability to search different levels of speech features while avoiding potential conflicts among them, during each mini-batch training, Choice Cell randomly samples one level of SEC, denoted as $X^{s}_{i}$, where $i\in[1,m,k]$, for calculation.
Secondly, for effective adjustment of TEC and to ensure that $X^{s}_{1}$ and $X^{s}_{m}$ do not contain target-related information, we detach $X^{s}_{i}$ before feeding it into Fusion Cells along with $X^{t}$.
Finally, following the previous fusion work, we offer a total of eight operations in the operation pool $\mathbb{O}$ of one Fusion Cell, and each operation takes two tensor inputs and yields a output:
$\small\mathrm{Zero}(X^{s}_{i},X^{t})=0$ (1)
$\small\mathrm{Sum}(X^{s}_{i},X^{t})=X^{s}_{i}+X^{t}$ (2)
$\mathrm{ConcatFC}(X^{s}_{i},$ $X^{t})=$ (4)
$\mathrm{ReLU}($ $\mathrm{Linear}(\mathrm{Concat}(X^{s}_{i},X^{t})))$
$\mathrm{where}\ H=\mathrm{Linear}(X^{t})$
[…]
2.3 Search Algorithm
In the Fusion Cell, we employ a differentiable search algorithm [5] to search the fusion operations:
where $\alpha\in\mathbb{R}^{|\mathbb{O}|}$ is a trainable weight that measures the importance of different operations, and $|\mathbb{O}|$ denotes the number of operations in $\mathbb{O}$. With the assistance of NAS, it is possible to avoid exhaustive training of all $3*8^{N_{c}}$ structures and find the optimal adjustment strategy through a single training process. Both $X^{s}_{k}$ and $X^{t}_{f}$ are subsequently used for emotional analysis.
3 Different modeling methods
To choose the most appropriate knowledge for extracting SEC, we conduct a comprehensive analysis of various pre-training modeling methods. The current modeling methods can be broadly categorized into two types: quantization-based reconstruction [6, 7, 8] and continuous frame-based reconstruction [24, 13]. Both approaches aim to understand the content of speech by reconstructing the masked portions. The key difference between these approaches lies in their representation space of speech frames. Formally, we represent the speech as $X=[x_{1},...,x_{T}]$, where $T$ is the number of frames. Given a set of indices $Z=[z_{1},...,z_{T^{\prime}}]$ corresponding to the frames that need to be masked, where $0\leq T^{\prime}\leq T$, the frames at the indices in $Z$ are replaced with a trainable code word $c$, resulting in $\tilde{X}=[...,x_{t},...,c,...]$, where $0\leq t<T$. The model then encodes $\tilde{X}$ to obtain its output $\hat{X}=[\hat{x}_{1},...,\hat{x}_{T}]$.
Figure 3: Unified training framework for different modeling.
Table 2: Performance comparison with works using text feature.
Speech
Feature
Text
Feature
Model
IEMOCAP
Leave-one-session
MFCC - BLSTM[25] 61.1 / 64.3
- ASR+BERT BLSTM[25] 71.8 / 71.9
MFCC ASR+BERT SAWC[25] 76.8 / 76.6
SEC TEC MSFN 74.0 / 71.9
@@ 4 Datasets
Interactive Emotional Dyadic Motion Capture (IEMOCAP) [26] dataset is a widely adopted benchmark for evaluating emotion recognition models. Following prior research, we focus on four emotions: angry, sad, happy, and neutral. To ensure a robust evaluation of our method, we adopted the 5-fold leave-one-session-out and 10-fold leave-one-speaker-out validation strategies to generate results.
@@
5 Experimental Setup & Results
5.1 Main setup
@@
In MFSN, we utilize pre-trained knowledge based on continuous modeling from Data2vec [13] to extract SEC from speech. For TEC, we retain prior knowledge of pronunciation units and capture it using fine-tuned Wav2vec2 [7]. Additionally, we employed an SGD optimizer with a learning rate of 5 to optimize the operation weight $\alpha$, as described in Section 2.3. For the convenience of analysis, we employ a single Fusion Cell. When dealing with MFSN dual-stream inputs, we double the dimension of the encoding hidden features for both spectrograms and MFCCs. These features are evenly distributed to each stream for co-attention calculations.
5.2 Comparison with existing works
Following prior works, we conducted performance tests of MFSN on both the IEMOCAP and SAVEE datasets. As depicted in Table 1, in our experiments on the IEMOCAP dataset, MFSN attains the highest Unweighted Accuracy (UA) score in the Leave-one-session cross-validation strategy, demonstrating competitive Weighted Accuracy (WA) scores even in the presence of class imbalance. Furthermore, within the Leave-one-speaker cross-validation strategy, MFSN consistently achieves the top UA and WA scores. When applied to the broader range of categories in the SAVEE dataset, MFSN exhibits a significant advantage, surpassing the sota in the 10-fold cross-validation.
However, as shown in Table 2, there is still a gap between TEC and BERT. Although MFSN surpass the performance of ASR+BERT, it still lags behind the SAWC [25].
5.3 Comparison between different modeling methods
5.3.1 Unified model configuration
@@ V A D UA(%) WA(%) V A D UA(%) WA(%) UA(%) WA(%)
As shown in Figure 3, to mitigate the impact caused by variations in engineering implementations, we employ a unified framework for different modeling methods. The specific parameters of the network are shown in Table 3. In the quantization modeling method, we imitate Wav2vec2 [7] and assign $B$ codebooks to all frame features, each containing $W$ codewords, resulting in a total of $N=W^{B}$ clustering centers. The label vector $Y$ is composed of the clustering centers corresponding to $X[Z]$ and the sampled negative examples. In the continuous modeling method, we imitate Data2vec [13] and use the mean value of the outputs of the last $L$ encoder layers as the $Y$. They were trained separately using BCE loss and MSE loss for reconstruction, with a data mask rate set at 65%.
5.3.2 Results
To fully compare the capabilities of various modeling methods in capturing SEC, we not only test the models’ capability in discrete emotion recognition but also in dimensional emotion analysis. In Table 4, we report the MSE metrics of Valence (V), Activation (A), Dominance (D), and report UA, WA.
As demonstrated by the performance in V, A, and D, in dimensional emotion analysis, knowledge acquired through quantization modeling exhibits notable limitations. The performance significantly improves as the number of quantization units, $N$, increases. Only when $N$ is large enough can this method approach the performance of continuous modeling. In the continuous modeling, the variation of the encoding layers $L$ as reconstructed labels does not noticeably affect performance.
@@ 5.3.2 Results
The above experimental results demonstrate that, whether in discrete emotion recognition or dimensional emotion analysis, the pre-trained knowledge derived from continuous modeling proves to be a superior choice for extracting SEC.
@@
5.4 Comparison with different configurations
We further analyze the importance of each module in MFSN by evaluating various model configurations using the Leave-one-session strategy on IEMOCAP. As shown in Table 5, consistent with the findings in Section 5.3.2, the pre-trained knowledge based on continuous modeling in Data2vec better captures SEC. However, simply incorporating TEC without adjustment, resulting in no performance improvement and even a slight decrease in the WA metric. Through the adjustment strategy search by MFSN, the model leverages SEC to mitigate bias in TEC, leading to a further improvement in model performance.
5.5 Visualization of Search Results
As shown in Figure 4, in various data partition scenarios, the Choice Cell and Fusion Cell automatically select optimal speech features and fusion operations to adjust TEC. To further demonstrate the performance of MFSN on imbalanced datasets, we present the performance confusion matrix of MFSN compared to the Co-attention method using only Wav2vec2 as the TEC Encoder, as shown in Figure 4.
Figure 4: The visualization of adjustment strategy search results for the Leave-one-session strategy. Here, $1$, $m$, and $k$ represent three levels of features. Red color indicates the best path.